OpenStudy (anonymous):

Solve the equation for g. g+25=c total cost for the ingredients: $29.40

OpenStudy (texaschic101):

does c stand for the cost of ingredients ?

OpenStudy (anonymous):

c = total cost

OpenStudy (texaschic101):

g + 25 = c c = 29.40 now just sub g + 25 = 29.40 --- subtract 25 from both sides g = 29.40 - 25 g = 4.40 <=== here you go

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

thank you so much!!!!!!

OpenStudy (texaschic101):

no problem :)